OpenStudy (anonymous):

3^x -2 = 27 Choose... x=0 x = 2/3 x=-6 x=5 x=2 x=-9

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

3^5 -2 =3^3=27

OpenStudy (anonymous):

thanks

OpenStudy (compassionate):

Myko, don't give out the answer.

OpenStudy (anonymous):